OpenStudy (anonymous):

f(x) = sin x - 6cos 3x when F(π/3) = 1

OpenStudy (anonymous):

\[f\left( x \right) = \sin x-6\cos 3x\]when \[f \left( \pi \div3 \right) = 1\] Hopefully this is a better way to explain my problem.

OpenStudy (anonymous):

\[\sin (x)-6 \cos (3 x)=1 \]when x = pi/2